Prove that 𝑓(𝑥) = 𝑥 2 + 2𝑥 is not injective


1
Expert's answer
2021-04-29T16:47:12-0400

Given,

"f(x)=x^2+2x"


"x^2+2x-y=0"


The solution of the above equation is-


"x=\dfrac{-2\pm\sqrt{4+4y}}{2}"


"x=-1\pm\sqrt{1+y}"


which is defined in R for "y\u2265\u22121." This can be used to separate the domain of f(x) into two intervals. For "x\in[1,∞" ) and "x\in(−∞,1]" , the function f(x) is not injective (and invertible). The domain that contains 2 is of course "[1,\u221e)."